# Theory of intervalley-coherent AFM order and topological superconductivity in tWSe2

### v1

@@ IV.1 Spin-fluctuation exchange as pairing mechanism in tWSe2
This procedure segregates spin fluctuations from other possible fluctuations captured by the FRG. The resulting RPA-renormalized particle-hole susceptibility $\hat{\chi}^{\text{RPA}}_{\uparrow\downarrow\uparrow\downarrow}(\bm{q})$ at scale $\Lambda=0.1\,\mathrm{meV}$ is shown in Fig. 3 (a). As expected from the Fermi surface sheets in the different spin-valley sectors, c.f. Fig. 1 (c), the dominant ordering vector $\bm{Q}_{C}$ locks to $K^{\nu}$ at the HOVHS (blue circle) and successively shifts towards incommensurate momenta on the $K^{\nu}$-$M$ line when lowering the external displacement field $E_{z}$. In this regime, the Fermi surface consists of triangular pockets around $K^{\nu}$ featuring three inequivalent van-Hove points associated to the three neighboring $M$-points, see Fig. 1 (c). Scattering between the van-Hove points enhances the particle-hole susceptibility leading to a pronounced commensurate-to-incommensurate transition of the IVC-AFM state along the van-Hove line, which is in good agreement with the interacting particle-hole susceptibility obtained from the full-fledged FRG treatment shown in Fig. 2 (b). To address pairing instabilities originating from fluctuations around the (incommensurate) IVC-AFM order that are present in the paramagnetic normal-state, we construct an effective pairing vertex $\Gamma^{P}(\bm{k}_{P},\bm{k}^{\prime}_{P})$ from the RPA renormalized interaction in the exchange channel Takimoto et al. (2004); Fischer et al. (2021) and analyze the symmetry of the leading superconducting order parameter by solving a linearized gap equation in orbital space, see SM SM for technical details. As shown in Fig. 3 (b), we find a superconducting state with mixed-parity that transforms in the two-dimensional $E$ irreducible representation ($d/p$-wave) in leading order for all values of the displacement field, exemplified for $E_{z}=25\,\mathrm{meV}$ (blue square). The superconducting order parameter is enhanced on the triangular Fermi pockets around $K^{\nu}$ that are exclusively $T/H_{2}$ polarized. Meanwhile, the amplitude of the SC order parameter is significantly suppressed if pairing involves Bloch states associated to the Wannier orbital $H_{1}$. This manifests in a vanishing SC amplitude on the larger Fermi surface arcs that connect neighboring $K^{\nu}$ points, c.f. Fig. 1 (c). Therefore, we expect that the superconducting state is effectively gapless over large parts of the Fermi surface leading to nodal signatures in spectroscopy measurements. The emergence of superconductivity from spin-fluctuation exchange can be understood by the real-space structure of the effective RPA pairing vertex $\Gamma^{P}(\bm{k}_{P},\bm{k}^{\prime}_{P})$ that features attractive components on bonds connecting neighboring $(T,H_{2})$ orbitals such to overcome the initially repulsive Coulomb interaction. As the phenomenology of the IVC-AFM state and the symmetry of the superconducting order parameter accurately agrees with the FRG results, we argue that spin fluctuations represent the pairing glue for electron-mediated superconductivity observed in tWSe2 along the van-Hove line.
Figure 4: Asymmetry of particle-hole instabilities due to Fermi surface broadening induced nesting. (a) Critical (relative) interaction scale $\alpha_{c}$ of the RPA renormalized interaction for different scales $\Lambda$ along the van-Hove line. The leading momentum transfer $\bm{Q}_{C}$ along the high symmetry path $K^{\nu}$-$M$ is encoded by the colorbar shared among panels (a,b). The minimal value of $\alpha_{c}$ for each curve is indicated by the gray dashed line showing a clear asymmetry of particle-hole instabilities with respect to the position of the HOVHS (blue circle). The lower panel shows the ($\Lambda$-broadened) Fermi surface sheets of the spin-valley polarized sub-sectors, i.e., all eigenenergies that are within the energy shell $[E_{F}-\Lambda,E_{F}+\Lambda]$ around the Fermi energy $E_{F}$. (b) Same information as in panel (a), taken for the constant displacement field cut $E_{z}=29\,\mathrm{meV}$ as indicated by the grey arrows in Fig. 2 (a). The grey solid line indicates the DOS as the density $n/n_{0}$ is varied across the van-Hove line. For low scales, the minimal value of $\alpha_{c}$ (grey dashed line) is encountered on the van-Hove line (green arrow) and the leading momentum transfer $\bm{Q}_{c}$ is displaced from $K^{\nu}$. At larger scales, $\alpha^{\text{min}}_{c}$ is shifted to larger hole fillings and $\bm{Q}_{c}$ locks back to $K^{\nu}$.
IV.2 Asymmetry of particle-hole instabilities by Fermi surface broadening induced nesting
An interesting observation of recent experiments Xia et al. (2024); Guo et al. (2024) concerns the apparent asymmetry of instabilities with respect to the displacement field tunable VHS that also manifests in the FRG phase diagram, i.e. (i) the maximal transition temperature of the IVC-AFM phase is encountered for values of the displacement field below the position of the HOVHS and (ii) instabilities are enhanced for $n<n_{\mathrm{VHS}}$, but diminish for $n>n_{\mathrm{VHS}}$. Here, we demonstrate that in the weak-to-moderate interacting regime, this asymmetry is related to scale-dependent nesting properties of the Fermi surface Mai et al. (2022); Šimkovic IV et al. (2022) that lead to commensurability locking of particle-hole instabilities away from the van-Hove line. To analyze the aforementioned asymmetric behavior, we again concentrate on a single-channel RPA resummation in the crossed particle-hole channel and calculate the critical (relative) interaction strength $\alpha_{c}$ defined by the divergence of the RPA renormalized interaction $C^{\mathrm{RPA}}_{U}(\bm{q})$ [cf. Eq. 4], i.e., $\det[\mathbb{1}-\alpha_{c}\hat{U}\hat{\chi}^{0}(\bm{q})]=0$. Figure 4 (a) displays the critical value $\alpha_{c}$ and the leading momentum transfer $\bm{Q}_{C}$ of the (incommensurate) IVC-AFM state along the van-Hove line for different RG scales $\Lambda$. The latter enters the RPA theory as temperature in the Fermi-Dirac statistics and hence determines the smearing of states in the vicinity of the Fermi energy, see SM SM for technical details. At low scale $\Lambda=0.1\,\mathrm{meV}$, the minimal critical interaction $\alpha_{c}^{\text{min}}$ (gray dashed line) is found at the HOVHS (blue circle), where the leading momentum transfer of the associated IVC-AFM state locks to $\bm{Q}_{C}=K^{\nu}$. This finding is consistent with the enhanced DOS at the HOVHS and the fact that the vector $\bm{q}=K^{\nu}$ perfectly connects Fermi surface sheets (FS) of opposite spin-valley polarization. To this end, we show the $\Lambda$-broadened FS, i.e., all eigenenergies that reside within the energy shell $[E_{F}-\Lambda,E_{F}+\Lambda]$ around the Fermi energy $E_{F}$, in the lower panel of Fig. 4 (a). Lowering the displacement field continuously displaces $\bm{Q}_{C}$ towards the $M$-point and increases the critical interaction strength $\alpha_{c}$ as scattering between the triangular FS pockets around $K^{\nu}$ is suppressed due to the relative orientation mismatch of the van-Hove points in different spin-valley sectors. At larger scales $\Lambda=\{0.5,1.0\}\,\mathrm{meV}$, however, the suppression of particle-hole instabilities for $E_{z}<E_{z}^{c}$ and the associated shift of $\bm{Q}_{C}$ towards the $M$-point are counteracted by the induced smearing of the triangular FS pockets formed by the three van-Hove points. Lowering the displacement field along the van-Hove line (left panel) increases the size of the FS pockets around $K^{\nu}$ such that scattering with vector $\bm{Q}_{C}=K^{\nu}$ between FS sheets of opposite spin-valley polarization is (re)enhanced at larger scales $\Lambda$. As a consequence of the higher scale $\Lambda$, the minimal critical value $\alpha_{c}^{\mathrm{min}}$ moves to lower values of the displacement field along the van-Hove line. This causes an asymmetry of interaction-induced particle-hole instabilities with respect to the HOVHS.
A similar asymmetry is encountered when varying the hole density $n/n_{0}$ for constant values of the displacement field $E_{z}<E_{z}^{c}$ as demonstrated in Fig. 4 (b). To show this, we focus on $E_{z}=29\,\mathrm{meV}$ (gray dash), where the asymmetry is maximal in the FRG phase diagram as demonstrated in Fig. 2 (a). In this regime, the Fermi surface at the VHS ($n_{\mathrm{VHS}}\sim-1.3$, green arrow) consists of three van-Hove points that are displaced from $K^{\nu}$ such that the leading momentum transfer $\bm{Q}_{C}$ is shifted towards the $M$-point. Decreasing the hole density lifts the van-Hove points and increases the size of the (open) FS pockets, similar to the observed phenomenology when varying the displacement field along the van-Hove line. For larger scales (temperatures), the broadening of the FS pockets (re)enhances scattering processes with momentum transfer $\bm{Q}_{C}=K^{\nu}$ and the minimal critical interaction strength $\alpha_{c}^{\text{min}}$ is moved to hole densities $n<n_{\mathrm{VHS}}$.
In conclusion, we argue that the minimal critical interaction $\alpha_{c}^{\text{min}}$ is shifted to lower displacement fields $E_{z}<E_{z}^{c}$ and lower densities $n<n_{\mathrm{VHS}}$ with respect to the position of the HOVHS and the displacement field tunable van-Hove line when particle-hole instabilities are encountered at non-zero temperature, or higher scales $\Lambda$. As the critical scale $\Lambda_{c}$ of particle-hole instabilities encountered in the FRG flow depends primarily on the strength of the Hubbard interactions $U_{X}$ in Eq. 1, it is an indicator for how strongly coupled the system under scrutiny is. Therefore, it is expected that the observed asymmetry is enhanced when increasing the strength of the interactions relative to the bandwidth. This can be achieved by tuning the twist angle (see Section VI) or altering the dielectric environment, i.e., changing the distance to metallic gates. We further note that the scale-dependent nesting properties of tWSe2 are tied to microscopic details of the fermiology of the system, particularly the formation of the triangular FS pockets around $K^{\nu}$ and the flat regions in momentum space originating from the HOVHS. This asymmetric occurrence of particle-hole instabilities and the associated commensurability locking may hence not be unique to tWSe2, but appear generically in systems with tunable FS pockets and HOVHS.
Figure 5: Influence of long-ranged density-density interactions on the low-energy phase diagram of $\theta=5.08^{\circ}$ tWSe2. (a) FRG phase diagram as function of holes per moiré unit cell $n/n_{0}$ and external displacement field $E_{z}$. The on-site IVC-AFM order parameter mixes with a spin-bond order (SBO) parameter between the $T$/$H_{2}$ orbitals that transform in the $A$-IR of $C_{3z}$. This raises the critical scale $\Lambda_{\mathrm{c}}$ of the density wave (DW) instabilities in the phase diagram and enhances the critical temperature of the superconducting phase. (b) Real-space structure of the particle-hole instabilities. The (magnetic) Wigner-Seitz cell is indicated by the (green) gray dashed line.
V Influence of long-ranged interactions: Formation of spin-bond order
@@
Besides the dominant Hubbard interaction terms acting on the triangular ($T$) and honeycomb sites ($H_{1,2}$) of the effective Wannier model defined in Eq. 1, the ratio between the Wannier function spread $\Omega\sim|\bm{a}_{M}|=3.74\,\mathrm{nm}$ and the distance to the metallic gates $\xi\sim 10\,\mathrm{nm}$ on either side of the sample suggests that longer-ranged density-density interactions impact the low-energy physics of $\theta=5.08^{\circ}$ tWSe2. Meanwhile, exchange and pair-hopping contributions between neighboring Wannier orbitals are negligible as expected from the exponential localization of the Wannier orbitals, see SM SM . Therefore, we next complement the Hubbard Hamiltonian Eq. 1 by density-density terms that are obtained when projecting the dual-gated Coulomb interaction to the Wannier basis:
@@
In the spirit of an ab-initio characterization of correlated phases in tWSe2, we consider the full real-space dependence of $V_{XX^{\prime}}^{\bm{R}\bm{R}^{\prime}}$ in this Section. The largest of these density-density terms are the nearest-neighbor honeycomb-triangular (honeycomb-honeycomb) density-density interactions $V_{TH}\approx 33\,\mathrm{meV}$ ($V_{HH}\approx 28\,\mathrm{meV}$). We demonstrate in the SM SM that truncating Eq. 5 after these terms already qualitatively accounts for the effects of longer-range interactions observed below.
@@
The corresponding FRG phase diagram for $\theta=5.08^{\circ}$ tWSe2 is shown in Fig. 5 (a). While the qualitative features of the phase diagram resemble those of the pristine Hubbard model Eq. 1, the phase boundaries and critical scales of the emergent particle-hole (particle-particle) instabilities are altered. The observed quantitative changes arise because the inter-site interaction causes non-local particle-hole pairs to get admixed to the local IVC-AFM order parameter. The nearest-neighbor repulsion between triangular ($T$) and honeycomb ($H_{2}$) orbitals supports the formation of a spin/valley-bond order (SBO)
@@
The SBO transforms within the trivial $A$-irreducible representation (IR) of the little group $C_{3z}$ at $K^{\nu}$ and is therefore symmetry-allowed to mix with the IVC-AFM state $\Delta^{\mathrm{AFM}}$ driven by the on-site Hubbard repulsion. In Fig. 5 (b), we present a sketch of the magnetization patterns in the IVC-AFM+SBO phase comprising an on-site $120^{\circ}$ spin configuration as well as the bond order component between $T$-$H_{2}$. Note that due to the $p_{\nu}$-orbital character of the $H_{2}$-orbital, the bonds feature a relative phase of $\omega=e^{2\pi i/3}$ between $C_{3z}$-related $H_{2}$-orbitals. The gain in free energy arises because the projection of the bare nearest-neighbor repulsion $V_{TH(HH)}$ to the crossed particle-hole channel lowers the energy of states with an SBO component of the order:
@@
see SM SM for details. The non-vanishing SBO components in the order parameter hence promote the formation of particle-hole pairs at larger critical scales $\Lambda_{c}$ as demonstrated in Fig. 5 (a) [compare with Fig. 2 (a)]. We therefore argue that the particle-hole instability experiences an effective anti-screening by virtue of long-ranged density-density type interactions. The coupling to the on-site IVC-AFM order that drives the chiral $p/d$-wave pairing instability causes the maximal critical temperature to rise to $\Lambda_{c}\sim 1\,\mathrm{K}$. Meanwhile, the size of the SC region in the phase diagram is decreased due to the competition with the IVC-AFM state.
Figure 6: Evolution of IVC-AFM domains and SC order in the low-energy phase diagram of tWSe2 in the intermediate twist angle regime $\theta=4^{\circ}\dots 5^{\circ}$. (a) FRG phase diagram at $\theta=4.5^{\circ}$ tWSe2 considering Hubbard-type interactions on the triangular ($T$) and honeycomb ($H_{1,2}$) orbitals. The blue circle indicates the position of the HOVHS. Regions of IVC-AFM order evolve to smaller absolute values of the displacement field $E_{z}$ and larger critical scales signaling the onset of correlated states at higher critical temperatures $\Lambda_{\mathrm{c}}\sim T_{\mathrm{c}}$. The maximal critical scale $\Lambda^{\text{max}}_{\mathrm{c}}$ features a pronounced asymmetry with respect to the HOVHS (blue circle) and the tunable van-Hove line (grey). (b) FRG phase diagram at $\theta=4.1^{\circ}$. Reaching the moderate-to-strongly interacting regime, the critical scales of IVC-AFM and chiral $d/p$-wave superconductivity are further enhanced. As a consequence, regions of SC order in the $E_{z}-n/n_{0}$ phase diagram are shrunk in size in favor of IVC-AFM ordered regions that move towards $n/n_{0}\sim 1$.
VI Evolution of the SC phase in the intermediate twist angle regime

### v2

@@ VI.1 Spin-fluctuation exchange as pairing mechanism in tWSe2
This procedure segregates spin fluctuations from other possible fluctuations captured by the FRG. The resulting RPA-renormalized particle-hole susceptibility $\hat{\chi}^{\text{RPA}}_{\uparrow\downarrow\uparrow\downarrow}(\bm{q})$ at scale $\Lambda=0.1\,\mathrm{meV}$ is shown in Fig.˜5 (a). As expected from the Fermi surface sheets in the different spin-valley sectors, c.f. Fig.˜1 (c), the dominant ordering vector $\bm{Q}_{C}$ locks to $K^{\nu}$ at the HOVHS (blue circle) and successively shifts towards incommensurate momenta on the $K^{\nu}$-$M$ line when lowering the external displacement field $E_{z}$. In this regime, the Fermi surface consists of triangular pockets around $K^{\nu}$ featuring three inequivalent van-Hove points associated to the three neighboring $M$-points, see Fig. 1 (c). Scattering between the van-Hove points enhances the particle-hole susceptibility leading to a pronounced commensurate-to-incommensurate transition of the IVC-AFM state along the van-Hove line, which is in good agreement with the interacting particle-hole susceptibility obtained from the full-fledged FRG treatment shown in Fig.˜2 (b). To address pairing instabilities originating from fluctuations around the (incommensurate) IVC-AFM order that are present in the paramagnetic normal-state, we construct an effective pairing vertex $\Gamma^{P}(\bm{k}_{P},\bm{k}^{\prime}_{P})$ from the RPA renormalized interaction in the exchange channel [102, 103] and analyze the symmetry of the leading superconducting order parameter by solving a linearized gap equation in orbital space, see SM [76] for technical details. As shown in Fig.˜5 (b), we find a superconducting state with mixed-parity that transforms in the two-dimensional $E$ irreducible representation ($d/p$-wave) in leading order for all values of the displacement field, exemplified for $E_{z}=25\,\mathrm{meV}$ (blue square). The emergence of superconductivity from spin-fluctuation exchange can be understood by the real-space structure of the effective RPA pairing vertex $\Gamma^{P}(\bm{k}_{P},\bm{k}^{\prime}_{P})$ that features attractive components on bonds connecting neighboring $(T,H_{2})$ orbitals such to overcome the initially repulsive Coulomb interaction. As the phenomenology of the IVC-AFM state and the symmetry of the superconducting order parameter accurately agrees with the FRG results, we argue that spin fluctuations represent the pairing glue for electron-mediated superconductivity observed in tWSe2 along the van-Hove line.
VI.2 Orbital-selective superconductivity
Next, we discuss the spectral and topological properties within the ordered SC phase as summarized in Fig. 6. The amplitude of the superconducting order parameter defined as
$|\Delta(\bm{k})|=\sqrt{\frac{1}{2}\text{Tr}\left[\hat{\Delta}^{\dagger}(\bm{k})\hat{\Delta}(\bm{k})\right]}\,.$ (12)
is shown in Fig. 6 (b) and is enhanced on the triangular Fermi pockets around $K^{\nu}$ that feature the largest momentum-resolved DOS and are exclusively $T/H_{2}$-polarized, c.f. Fig. 1 (c). As the SC order parameter only has weight on bonds including the $T/H_{2}$-orbital, whereas $\Delta_{H_{1}X}(\bm{k})=\Delta_{XH_{1}}(\bm{k})=0$ for $X\in\{T,H_{1},H_{2}\}$, the amplitude is significantly suppressed on those Fermi surface sheets that have non-vanishing $H_{1}$-polarization. This manifests in a vanishing SC amplitude on the dispersive Fermi surface sheets that connect neighboring $K^{\nu}$ points as shown in Fig. 6 (b). Therefore, the orbital-selective superconducting state is effectively gapless over large parts of the Fermi surface leading to nodal signatures in spectroscopy measurements. To underline this statement, we calculate the local density of states (LDOS) in the SC phase as shown in Fig. 6 (a). While the $H_{2}$-orbital that is centered at the MX stacking regions of the moiré unit cell features a full SC gap visible by the ‘U’-shape of the LDOS, the MM and XM stacking regions show nodal ‘V’-shape characteristics. These nodal signatures are associated to the orbital-selectivity of the SC order parameter exhibiting strong suppression of the pairing amplitude on the dispersive Fermi arcs with mixed $T/H_{1}$ polarization. We therefore claim that the SC order parameter, albeit being fully gapped around $K^{\nu}$, features nodal signatures in spectroscopy measurements.
The chiral superconducting state of $\theta=5.08^{\circ}$ tWSe2 belong to the A Cartan class and possesses a $\mathbb{Z}$ topological invariant as outlined in Section III.3. Here, we explicitly calculate the non-abelian Berry curvature
$\bm{\Omega}(\bm{k})$ ${}=-\text{Im}\big\{\log\det_{\mu\nu}\langle\nabla_{\bm{k}}u_{\mu}(\bm{k})|\times|\nabla_{\bm{k}}u_{\nu}(\bm{k})\rangle\big\}$ (13)
where $u_{\mu}(\bm{k})$ is the (Nambu) Bloch wave function at crystal momentum $\bm{k}$ and the band index $\mu$ comprises all occupied states of the Nambu Hamiltonian. The Chern number is then obtained by $C=1/(2\pi)\sum_{\bm{k}}\Omega_{z}(\bm{k})$. We show the distribution of Berry curvature within the moiré BZ in Fig. 6 (c). Berry curvature predominantly accumulates around $K^{\nu}$, where the SC order parameter fully gaps the system and along the dispersive Fermi arcs, along which the SC amplitude is drastically suppressed.
Figure 6: Spectral and topological properties of the chiral $d/p$-wave SC order parameter. (a) Local density of states (LDOS) in the superconducting phase resolved for the Wannier orbitals $X\in\{T,H_{1},H_{2}\}$ that reside at the $M$, XM and MX stacking regions of the moiré unit cell respectively. (b) Momentum-resolved amplitde of the superconducting order parameter in the moiré Brillouin zone. (c) Non-abelian Berry curvature $\Omega(\bm{k})$ over the occupied Nambu bands of the Bogoliubov de-Gennes Hamiltonian. The chiral $d/p$-wave SC order parameter features a non-vanishing Chern number $C=2$.
VII Influence of long-ranged interactions on the low-energy phase diagram
@@
Besides the dominant Hubbard interaction terms acting on the triangular ($T$) and honeycomb sites ($H_{1,2}$) of the effective Wannier model defined in Eq.˜2, the ratio between the Wannier function spread $\Omega\sim|\bm{a}_{M}|=3.74\,\mathrm{nm}$ and the distance to the metallic gates $\xi\sim 10\,\mathrm{nm}$ on either side of the sample suggests that longer-ranged density-density interactions impact the low-energy physics of $\theta=5.08^{\circ}$ tWSe2. Meanwhile, exchange and pair-hopping contributions between neighboring Wannier orbitals are negligible as expected from the exponential localization of the Wannier orbitals, see SM [76]. Therefore, we next complement the Hubbard Hamiltonian Eq.˜2 by density-density terms that are obtained when projecting the dual-gated Coulomb interaction to the Wannier basis:
@@
In the spirit of an ab-initio characterization of correlated phases in tWSe2, we consider the full real-space dependence of $V_{XX^{\prime}}^{\bm{R}\bm{R}^{\prime}}$ in this Section. The largest of these density-density terms are the nearest-neighbor honeycomb-triangular (honeycomb-honeycomb) density-density interactions $V_{TH}\approx 33\,\mathrm{meV}$ ($V_{HH}\approx 28\,\mathrm{meV}$).
VII.1 Influence of long-ranged density-density interactions on the position of the van-Hove line and the layer-polarization regime.
Before turning the discussion to the impact of long-ranged Coulomb interactions on the IVC-AFM order and superconducting regions in the phase diagram of $\theta=5.08^{\circ}$ tWSe2, we first investigate the effect of symmetry-unbroken moiré band renormalizations as function of external displacement field and filling. Hartree-induced band renormalizations have been shown to strongly renormalize the low-energy flat bands in twisted multilayer graphenes [104, 105, 106, 107, 108, 109, 110, 111] leading to the pinning of the van-Hove singularities of the flat moiré bands to the Fermi level such to enhance the ordering tendency of the system over extended filling regions. Here, we analyze the impact of symmetry-unbroken Hartree-Fock induced band renormalization by starting from the Hamiltonian $H^{V}$ defined in Eq. (14) and including the (static) self-energy $\Sigma(\bm{k})$ within the FRG formalism. Therefore, we effectively account for all Hartree-Fock type diagrams and their renormalization of the single-particle bandstructure, see Supplementary Material for details. To disentangle the role of interaction-induced band renormalizations and the formation of electronic order, we neglect the flow of the two-particle interaction vertex $\Gamma^{(4)}$ ($\Sigma$-flow) and study the impact of symmetry-unbroken self-energy corrections on the DOS in the phase diagram of displacement field and filling as showin in Fig. 7.
First and foremost the van-Hove line is more strongly coupled to the layer-polarized regime (black dashed line), in which the system is exclusively polarized on the $T/H_{1}$-orbital leading to a sudden suppression of the DOS. This pinning of the van-Hove line with maximal DOS and the layer-polarization line is facilitated by an additional band flattening near the $K,K^{\prime}$ points of the moiré BZ in the respective density regime, see Supplementary Material [76] for plots of the renormalized bandstructure as function of external displacement field and hole filling. Meanwhile, the fermiology along the van-Hove line is not altered by the symmetry-unbroken Hartree-Fock corrections as demonstrated in the upper panel of Fig. 7 (a) that show the Fermi surface on the van-Hove line for the same displacement fields as the normal-state DOS shown in Fig. 1. In the normal-state (black dashed line), the critical displacement field $E_{Z}^{\text{lp}}$ at which the system enters the ’layer-polarization’ regime (grey dots) and only the Wannier orbitals $T,H_{1}$ contribute to the Fermi surface is a strictly linear function of the external displacement field as expected from the interlayer potential difference $E_{z}$ encoded in the multi-orbital Wannier Hamiltonian Eq. (2). In the presence of long-ranged density-density interaction, the critical displacement fields picks up a non-linear dependence $E_{z}^{\text{lp}}\sim[n/n_{0}]^{2}$ at small densities $1<n/n_{0}<0$, in well agreement with experimental findings [39, 37, 38].
Figure 7: Symmetry-unbroken band renormalizations in the presence of long-ranged density-density interactions. The DOS is obtained at each point in the $(E_{z},n)$ phase diagram by a FRG flow of the self-energy. The critical displacement field $E_{Z}^{\text{lp}}$ at which the system enters the ’layer-polarization’ regime (black dots) and only the Wannier orbitals $T,H_{1}$ contribute to the Fermi surface picks up a non-linear dependence $E_{z}^{\text{lp}}\sim[n/n_{0}]^{2}$. To see this, we plot the layer-polarization line (grey dots) in the uncorrelated normal-state from Fig. 1 (c) as a reference. The fermiology along the tunable van-Hove line (inset) resembles the uncorrelated normal-state.
VII.2 FRG phase diagram: Formation of spin-bond order
@@
As the symmetry-unbroken band renormalizations from long-ranged density-density interactions only yield minor quantitative corrections to the position of the DOS in the phase diagram of displacement field and filling, we proceed by studying the FRG phase diagram for $\theta=5.08^{\circ}$ tWSe2 as shown in Fig.˜8 (a). While the qualitative features of the phase diagram resemble those of the pristine Hubbard model Eq.˜2, the phase boundaries and critical scales of the emergent particle-hole (particle-particle) instabilities are altered. The observed quantitative changes arise because the inter-site interaction causes non-local particle-hole pairs to get admixed to the local IVC-AFM order parameter. The nearest-neighbor repulsion between triangular ($T$) and honeycomb ($H_{2}$) orbitals supports the formation of a spin/valley-bond order (SBO)
@@
The SBO transforms within the trivial $A$-irreducible representation (IR) of the little group $C_{3z}$ at $K^{\nu}$ and is therefore symmetry-allowed to mix with the IVC-AFM state $\Delta^{\mathrm{AFM}}$ driven by the on-site Hubbard repulsion. In Fig.˜8 (b), we present a sketch of the magnetization patterns in the IVC-AFM+SBO phase comprising an on-site $120^{\circ}$ spin configuration as well as the bond order component between $T$-$H_{2}$. Note that due to the $p_{\nu}$-orbital character of the $H_{2}$-orbital, the bonds feature a relative phase of $\omega=e^{2\pi i/3}$ between $C_{3z}$-related $H_{2}$-orbitals. The gain in free energy arises because the projection of the bare nearest-neighbor repulsion $V_{TH(HH)}$ to the crossed particle-hole channel lowers the energy of states with an SBO component of the order:
@@
see SM [76] for details. The non-vanishing SBO components in the order parameter hence promote the formation of particle-hole pairs at larger critical scales $\Lambda_{c}$ as demonstrated in Fig.˜8 (a) [compare with Fig.˜2 (a)]. We therefore argue that the particle-hole instability experiences an effective anti-screening by virtue of long-ranged density-density type interactions. The coupling to the on-site IVC-AFM order that drives the chiral $p/d$-wave pairing instability causes the maximal critical temperature to rise to $\Lambda_{c}\sim 1\,\mathrm{K}$. Meanwhile, the size of the SC region in the phase diagram is decreased due to the competition with the IVC-AFM state.
[…]
VIII Evolution of the IVC-AFM order and superconductivity across twist angles